OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = 3x + 4, what is f(g(x)) ? is it 12x + 15? If f(x) = 4x – 1 and g(x) = 3x + 4, what is f(g(x)) ? is it 12x + 15? @Mathematics

hero (hero):

f(g(x)) = 4(3x+4)-1 = 12x+16 -1 = 12x+15 Correct

OpenStudy (fools101):

oh i know a place u can check if that right! www.mathway.com

OpenStudy (fools101):

hope this helps!

hero (hero):

Wolfram Alpha is light years better than what mathway can ever even dream to be.

OpenStudy (fools101):

oh kk hero for got about that one too but i use both : )